Solve the problem. Please show all of your work.
A college has a stairway in the business wing that has steps that are 8 inches high and 9.2 inches across. What is the slope of the stairs?

Answer by stanbon(75887)

A college has a stairway in the business wing that has steps that are 8 inches high and 9.2 inches across. What is the slope of the stairs?
----
slope = rise/run = 8/9.2 = 0.8696
=================================
Cheers,
Stan H.